# Aladin Djuhera1, Amin Seffo1, Vlad C. Andrei1, Holger Boche1, and Walid Saad2

### v2

@@ IV-B3 Path 3 - Extreme Case
Finally, we investigate an extreme case in Fig. 5, where a trajectory with threshold $G=0.6$ is requested for a very short path between $n_{\text{start}}=(1.5,-1)$ to $n_{\text{goal}}=(-1.5,-1)$. In this example, both A* and N-WA* maintain a short and roughly straight line from start to goal, where N-WA* biases toward the shortest path and thus cannot optimize the trade-off between optimal path gain and trajectory length. However, all SCoTT and DP paths take a semicircle-shaped detour into higher path gain areas with average gains of 0.69 for DP-WA*, 0.62 for SCoTT, and 0.66 for SCoTT-
Figure 3: Results for path planning with $G$=0.7. SCoTT immediately moves toward the high gain blue area and overlaps with the optimal DP-WA* therein.
Figure 4: Results for path planning with $G$=0.4. All wireless-aware approaches, including N-WA*, avoid the shortest path and instead take a detour.
Figure 5: Results for path planning with $G$=0.6. N-WA* biases toward the shortest path while SCoTT and DP flavors effectively optimize for $G$.
DP-WA*. Similarly as before, SCoTT further reduces the total path length while SCoTT-DP-WA* significantly reduces the computation time by 56% from 76.97s to 33.76s compared to DP-WA*. In this example, the threshold $G=0.6$ is specifically chosen to be higher in order to motivate such a stark detour from the shortest path, which can only be achieved by explicitly wireless-aware approaches such as DP-WA*, SCoTT, and SCoTT-DP-WA*.
@@
Path 1: Across the Room Path 2: Wall to Wall Path 3: Extreme Case
SCoTT Llama-4 Avg. Path Gain 0.72 0.41 0.62
@@
SCoTT Llama-3 Avg. Path Gain 0.71 0.42 0.61
@@
SCoTT SmolVLM Avg. Path Gain 0.78 0.49 0.67
@@
SCoTT Granite Avg. Path Gain 0.75 0.46 0.68
@@
IV-C Model Variations.
We ablated SCoTT’s generative model across four different open-source VLMs of varying size: Llama-4-Scout-17B [11], Llama-3.2-11B-Vision [12], SmolVLM [13], and Granite-Vision-3.2-2B [14]. All models were run locally on our compute cluster. Table II reports runtime and performance metrics for each model, averaged over $N=10$ runs. We also report a success ratio, defined as the fraction of runs that produced valid paths, i.e., those respecting the average path gain constraints and successfully avoiding obstacles in the DT.
In general, the larger models, Llama-4 and Llama-3, achieve the best balance between average path gain and total distance, with Llama-4 outperforming Llama-3 in both speed and path length. Surprisingly, the smaller 2B models, SmolVLM and Granite, remain highly competitive. However, these compact variants tend to favor higher gains at the expense of slightly longer routes, reflecting their bias toward blue- and green-rich regions during the visual reasoning step. This behavior increases their overall runtime compared to the Llama models, as more waypoints are generated in areas with higher path gains, albeit only marginally. In our experiments, nearly all models achieved a 100% success ratio, except SmolVLM, which occasionally failed on Paths 1 and 2, missing obstacles. These results demonstrate that SCoTT’s prompting strategy is robust across different model architectures and that it can leverage lightweight models effectively at lower inference cost, which is particularly important for real-world deployments.
@@ V System Integration and Practical Considerations
Figure 6: End-to-end integration of SCoTT with DTs and ROS. The DT supplies spatial information to a wireless ray-tracer, which generates wireless heatmaps and path gain measurements. SCoTT then plans a wireless-aware path that ROS nodes execute within the DT environment.
V-A System Integration with ROS.
Fig. 6 shows how SCoTT integrates with our DT and a dedicated ROS module. In this setup, SCoTT plans a wireless-aware path using multi-modal data from the wireless ray-tracer and spatial information provided by the DT. The result is a sequence of 2D waypoints that guide the robot from its start to a goal position. These waypoints feed into a ROS-based control loop for real-time navigation. In general, navigation differs from path planning in that it covers the execution of the planned trajectory, incorporating localization, motion control, and sensor feedback. A snapshot of this process is shown in Fig. 7, where the environment is rendered as a 2D occupancy grid map built with Simultaneous Localization and Mapping (SLAM) in ROS. The red line traces the robot’s actual trajectory, closely matching SCoTT’s planned path in black. Small deviations arise from real-time corrections as the robot compensates for odometry noise and minor localization drift, common in SLAM-based navigation. We observe consistent results across all of our path planning examples from Sec. IV-B. This pipeline completes the full cycle: mapping the environment for obstacle avoidance, performing wireless-aware path planning with SCoTT, and executing the mission through the ROS control stack, ensuring realistic motion execution and map consistency during the mission.
@@
V-B Practical Considerations.
SCoTT is fundamentally an inference‐only framework, such that no fine‐tuning or additional training is required. In addition, our experiments showed comparable performance using both large (Llama-4-Scout-17B) and compact (Granite-Vision-3.2-2B) models, suggesting that SCoTT can run on edge or near‐edge hardware with modest latency and at low inference cost. Moreover, SCoTT’s data pipeline is mainly driven by the DT and wireless ray‐tracer, which generate the heatmap images and path gain JSON. These assets can be fed into a lightweight RAG layer, which adds negligible overhead as data is handled on a per-need basis in each subtask. This model‐agnostic, modular design makes SCoTT both generalizable and practical for real‐time, 6G‐enabled robotic navigation.
@@ VI Conclusion
In this paper, we proposed SCoTT, a wireless‐aware path‐planning framework that leverages multi‐modal VLMs to jointly optimize average wireless path gains and trajectory length. By breaking a single planning query into a sequence of focused CoT subtasks, each guided by an explicit reasoning strategy and workflow, SCoTT overcomes the hallucination and context‐window limitations of conventional language-based planners and poses an alternative to prohibitive classical planners. Through extensive simulations and ablation studies on three challenging scenarios and four different VLM models, we demonstrated that SCoTT achieves near‐optimal path gains within 2% of our cost-optimal reference baseline (DP‐WA*). We also showed that SCoTT can be used as a neuro-symbolic extension to classical algorithms, resulting in up to 62% faster downstream runtimes. Furthermore, our model ablations showed that even compact models yield valid routes, preparing SCoTT for on‐device deployment at low inference cost. Finally, we validated SCoTT’s practicality by integrating it end-to-end into a ROS-based navigation pipeline within a DT, demonstrating seamless integration and confirming its readiness for real-time, 6G-enabled robotic applications.

### v3

@@ IV-B3 Path 3 - Extreme Case
Finally, we investigate an extreme case in Fig. 5, where a trajectory with threshold $G=0.6$ is requested for a very short path between $n_{\text{start}}=(1.5,-1)$ and $n_{\text{goal}}=(-1.5,-1)$. In this example, both A* and N-WA* maintain a short and roughly straight line from start to goal, where N-WA* biases toward the shortest path and thus cannot optimize the trade-off between optimal path gain and trajectory length. However, all SCoTT and DP paths take a semicircle-shaped detour into higher path gain areas with average gains of 0.69 for DP-WA*, 0.62 for SCoTT, and 0.66 for SCoTT-
Figure 3: Results for path planning with GG=0.7: SCoTT immediately moves toward the high gain blue area and overlaps with the optimal DP-WA* therein.
Figure 4: Results for path planning with GG=0.4: All wireless-aware approaches, including N-WA*, avoid the shortest path and instead take a detour.
Figure 5: Results for path planning with GG=0.6: N-WA* biases toward the shortest path while SCoTT and DP flavors effectively optimize for $G$.
DP-WA*. Similarly, as before, SCoTT further reduces the total path length, while SCoTT-DP-WA* significantly reduces the computation time by 56% from 76.97s to 33.76s, compared to DP-WA*. In this example, the threshold $G=0.6$ is deliberately chosen to be higher in order to motivate such a stark detour from the shortest path, which can only be achieved by explicitly wireless-aware approaches such as DP-WA*, SCoTT, and SCoTT-DP-WA*.
@@
Path 1:
Across the Room
Path 2:
Wall to Wall
Path 3:
Extreme Case
SCoTT Llama-4
Avg. Path Gain 0.72 0.41 0.62
@@
SCoTT Llama-3
Avg. Path Gain 0.71 0.42 0.61
@@
SCoTT SmolVLM
Avg. Path Gain 0.78 0.49 0.67
@@
SCoTT Granite
Avg. Path Gain 0.75 0.46 0.68
@@
IV-C Model Variations
We ablated SCoTT’s generative model across four different open-source VLMs of varying sizes: Llama-4-Scout-17B [13], Llama-3.2-11B-Vision [14], SmolVLM [15], and Granite-Vision-3.2-2B [16]. All models were run locally on our compute cluster. Table II reports runtime and performance metrics for each model, averaged over $N=10$ runs. We also report a success ratio, defined as the fraction of runs that produced valid paths, i.e., those respecting the average path gain constraints and successfully avoiding obstacles in the DT.
In general, the larger models, Llama-4 and Llama-3, achieve the best balance between average path gain and total distance, with Llama-4 outperforming Llama-3 in both speed and path length. Surprisingly, the smaller models, SmolVLM and Granite, remain highly competitive. However, these compact variants tend to favor higher gains at the expense of slightly longer routes, reflecting their bias toward blue- and green-rich regions during the visual reasoning step. This behavior increases the overall runtime compared to the Llama models, as more waypoints are generated in areas with higher path gains, albeit only marginally. In our experiments, nearly all models achieved a 100% success ratio, except for SmolVLM, which occasionally failed on Paths 2 and 3 by missing obstacles. These results demonstrate that SCoTT’s prompting strategy is robust across different model architectures and that it can leverage lightweight models effectively at lower inference cost, which is particularly important for real-world deployments.
@@ V System Integration and Practical Considerations
Figure 6: End-to-end integration of SCoTT with DTs and ROS. The DT supplies spatial information to a wireless ray tracer, which generates wireless heatmaps and path gain measurements. SCoTT then plans a wireless-aware path that is executed via ROS nodes within the DT environment.
V-A System Integration with ROS
Fig. 6 shows how SCoTT integrates with our DT via dedicated ROS modules. In this setup, SCoTT plans a wireless-aware path using multi-modal data from the wireless ray tracer and spatial information provided by the DT. The result is a sequence of 2D waypoints that guide the robot from its start to a goal position. These waypoints feed into a ROS-based control loop for real-time navigation. In general, navigation differs from path planning in that it covers the actual execution of the planned trajectory, including localization, motion control, and sensor feedback. A snapshot of this process is shown in Fig. 7, where the environment is rendered as a 2D occupancy grid map built with Simultaneous Localization and Mapping (SLAM). The red line traces the robot’s actual trajectory, closely matching SCoTT’s planned path in black. Small deviations arise from real-time corrections as the robot compensates for odometry noise and minor localization drift, common in SLAM-based navigation. We observe consistent results across all the examples from Sec. IV-B. This pipeline completes the full cycle: mapping the environment for obstacle avoidance, performing wireless-aware path planning with SCoTT, and executing the mission through the ROS control stack, ensuring realistic motion execution and map consistency.
@@ V-A System Integration with ROS
V-B Practical Considerations
SCoTT is fundamentally an inference‐only framework, such that no fine‐tuning or additional training is required. In addition, our experiments showed comparable performance using both large (Llama-4-Scout-17B) and compact (Granite-Vision-3.2-2B) models, suggesting that SCoTT can run on edge or near‐edge hardware with modest latency and at low inference cost. Moreover, SCoTT’s data pipeline is mainly driven by the DT and wireless ray tracer, which generate the heatmap images and path gain JSON files. These assets can be fed into a lightweight RAG layer, which adds negligible overhead as data is handled on a per-need basis in each subtask. This model‐agnostic, modular design makes SCoTT both generalizable and practical for real‐time, 6G‐enabled robotic navigation.
@@ VI Conclusion
In this paper, we proposed SCoTT, a wireless‐aware path‐planning framework that leverages multi‐modal VLMs to jointly optimize average wireless path gains and trajectory lengths. By breaking a single planning query into a sequence of strategic CoT subtasks, each guided by an explicit reasoning strategy and workflow, SCoTT overcomes the hallucination and context‐window limitations of conventional language model-based planners and poses an alternative to prohibitive classical planners. Through extensive simulations and ablation studies on three challenging scenarios and four different VLMs, we demonstrated that SCoTT achieves near‐optimal path gains within 2% of our cost-optimal reference baseline (DP‐WA*). We also showed that SCoTT can be used as a neuro-symbolic extension to classical algorithms, resulting in up to 62% faster runtimes. Furthermore, our model ablations showed that even compact models yield valid routes, making SCoTT suitable for on-device deployment at low inference cost. Finally, we validated SCoTT’s practicality by integrating it end-to-end into a ROS-based navigation pipeline within a DT, demonstrating seamless integration and confirming its readiness for real-time, 6G-enabled robotic applications.